OpenStudy (anonymous):

How many solutions does the equation have ? 6-a=1/3(9a-18)

OpenStudy (luckycoins888):

6-a = (9a-18)/3 6-a = 3(3a-6)/3 simplified 3 by 3 then will get 6-a = 3a-6 12 = 4a a= 12/4 a = 3 One solution

OpenStudy (anonymous):

thanks homie

OpenStudy (luckycoins888):

Welcome m8

OpenStudy (anonymous):

aye can u help me on a couple more ?

OpenStudy (luckycoins888):

Shure

OpenStudy (anonymous):

ku, hold up

OpenStudy (luckycoins888):

Kk

OpenStudy (luckycoins888):

Why dont you start a new thread though XD